QUESTION 1

You still produce Coco cola beverages drinks. Currently, the Market price is $2.50 per 20-ounce can. You estimate that the price elasticity of demand is 2.7 (in absolute value). If there is a decrease in supply due to a $1.00 per can increase in costs, then you predict:

A relatively___________________change in quantity demanded and a relatively______________change in price.

QUESTION 2

Tom and you are stranded in the backwoods. You can trap beavers or hunt elk. Tom needs 60 minutes to trap a beaver and 90 minutes to kill an elk. You need 20 minutes to trap a beaver and 60 minutes to kill an elk. What should the two of you do?
Answer #1

larger

smaller

since elasticity of demand is 2.7, demand is elastic, it means buyers are responsive to changes in price. So decrease in supply results in larger decrease in quantity demanded and smaller increase in price.

-------

Tom's opportunity cost of trapping= 60/90= 0.67

Your opportunity cost of trapping= 20/60=0.33

Since you have lower opportunity cost of trapping, you should trap beaver and Tom should kill elk.